OpenStudy (anonymous):

What is the formula of the number series 1;4;16;64...

OpenStudy (anonymous):

Here a=1, r=2nd term/1st term=3rd term/2nd term=...=4/1=4 Hence it is a g.p series with the formula of the n-th term a*r^(n-1)=1*4^(n-1)=4^(n-1)

OpenStudy (anonymous):

@tellie

OpenStudy (anonymous):

where n=1,2,3,....

OpenStudy (anonymous):

Thank you SO much! Appreciate it!

OpenStudy (anonymous):

no problem.:)